The healthcare provider prescribes meperidine (Demerol) 1.5 mg/kg PO Q4H prn pain.

The pharmacy supplies the medication with a dosage strength of 50 mg/ 5mL. The child

weighs 22lb.

How many mL’s would the nurse administer? ____________mL

Please show work by using dimensional analysis ONLY
Answer #1

Prescribed meperidine= 1.5 mg/kg PO Q4H prn pain

PO means orally, Q4H means every 4 hours , prn =when required

Chlid's weight= 22 lb= 22/2.2 kg= 10 kg

As per weight of child required dose= 1.5X 10 = 15 mg PO Q4H prn

Available dose strength= 50 mg/5 ml

= 10 mg/ml

Hence for the ordered dose required ml= 15/10 = 1.5 ml